How many liters of pure antifreeze must be added to a 5 liters of a 10% antifreeze solution to obtain 20% antifreeze solution ?
Answer by richwmiller(17219)

1*x+0.1*5=0.2(5+x)
1*x+0.5=1+0.2x
1x-0.2x=1-0.5
0.8x=0.5
x=0.625
check
1*0.625+0.1*5=0.2(5+0.625)
0.625+0.5=0.2(5.625)
1.125=1.125
ok